OpenStudy (anonymous):

amal distributed 12 more questionnaires than Pat. Eli distributed 15 fewer questionnaires than Pat. Martine distributed the same number of questionnaires as Amal. What is the difference between the number of questionnaires distributed by amal and the number distributed by Eli?

OpenStudy (anonymous):

A=12+P E=P-15 to find the difference, A-E = 12+P -P+15 = 27

OpenStudy (anonymous):

there's no way you might of made an error right?

OpenStudy (anonymous):

u could double check my work. people make mistakes. :)

OpenStudy (anonymous):

No, I will accept it - your brilliant and I thank you from the bottom of my heart

OpenStudy (anonymous):

u r welcome..